OpenStudy (anonymous):

three of the four numbers have a sum of 22. Th average of the four numbers is 4. what is the fourth number

OpenStudy (kropot72):

Let the unknown number be x. Then we can write: \[\large \frac{22+x}{4}=4\ .......(1)\] Now just solve (1) to find x.

OpenStudy (nurali):

Three of four numbers have a sum of 22. a + b + c = 22 : If the average of the four numbers is 4, a+b+c+d/4= 4 Replace a + b + c with 22 22+d/4= 4 multiply both sides by 4 22 + d = 16 d = 16 - 22 d = -6 is the fourth number